What is the pH of a solution that is 0.49 M K2HPO4 and 0.056 M K3PO4? Use the Acid-Base Table. Please see Buffer pH for assistance.
Answer #1

From Henderson- Hasselbalch equation

pH = pKa + log [PO43-]/[HPO4-]

From literature

pKa = 12.7

Also given that,

[PO43-] = 0.056 M

and [HPO4-] = 0.49 M

pH = 12.7 + log (0.056/0.49) = 11.76